# DDAT: Diffusion Policies Enforcing Dynamically Admissible Robot Trajectories

### v1

@@ Abstract
Diffusion models excel at creating images and videos thanks to their multimodal generative capabilities. These same capabilities have made diffusion models increasingly popular in robotics research, where they are extensively used for generating robot motion. However, the stochastic nature of diffusion models is fundamentally at odds with the precise dynamical equations describing the feasible motion of robots. Hence, generating dynamically admissible robot trajectories is a challenge for diffusion models. To alleviate this issue, we introduce DDAT: Diffusion policies for Dynamically Admissible Trajectories to generate provably admissible trajectories of black-box robotic systems using diffusion models. A sequence of states is a dynamically admissible trajectory if each state of the sequence belongs to the reachable set of its predecessor by the robot’s equations of motion. To generate such trajectories, our diffusion policies project their predictions onto a dynamically admissible manifold during both training and inference to align the objective of the denoiser neural network with the dynamical admissibility constraint. Due to auto-regressive nature of such projections as well as the black-box nature of robot dynamics, characterization of such reachable sets will be challenging. We thus enforce admissibility by iteratively sampling a polytopic under-approximation of the reachable set of a state onto which we project its predicted successor, before iterating this process with the projected successor. By producing accurate trajectories, this projection eliminates the need for diffusion models to continually replan, enabling one-shot long-horizon trajectory planning. We demonstrate that our proposed framework generates higher quality dynamically admissible robot trajectories through extensive simulations on a quadcopter and various MuJoCo environments, along with real-world experiments on a Unitree GO1 and GO2. Our code is available at https://iconlab.negarmehr.com/DDAT/.
@@ I Introduction
To address the dynamic feasibility issue in diffusion planning from offline data, we introduce Diffusion policies for Dynamically Admissible Trajectories (DDAT). These models use a chronological trajectory projector capable of transforming a generated sequence of states into an admissible trajectory. On the contrary to work [22], and inspired by [10, 16, 48], the projections in DDAT occurs throughout and after inference instead of only after inference. Differing from works [10, 16, 48], these projections are also implemented during training of the diffusion models to align their training objectives with their inference phase, as suggested in work [5]. Furthermore, works [5, 10, 16, 48] only concern themselves with a single constraint, which is significantly less challenging to enforce than dynamical admissibility due to its autoregressive nature. Another advantage of generating feasible trajectories is that we do not need to sample and denoise a large batch of trajectories hoping to find an admissible one, which provides significant computational gain compared to methods like [8]. In summary, our main contributions in this work are as follows.
@@
    We generate dynamically admissible trajectories with diffusion models.
@@
The remainder of this work is organized as follows. In Section II, we provide a survey of related works. In Section III, we introduce our problem formulation along with our framework. In Section IV, we define several projectors to enforce the dynamical admissibility of trajectories. In Section V, we incorporate these projectors into diffusion models. In Section VI, we implement our proposed Diffusion policies for Dynamically Admissible Trajectories. We present our simulations results on a quadcopter, the Gym Hopper, Walker and HalfCheetah, and the Unitree GO2 quadruped in Section VII, while Section VIII demonstrates the hardware implementation of DDAT on a Unitree GO1. Finally, we conclude the paper in Section X.
@@ II-A2 Hard constraints
Contrasting the works presented above, we are interested in hard constraints to guarantee their satisfaction. The first hard constraints imposed on diffusion models were $[0,255]$ pixel cutoffs, which led to saturated images [28]. To mitigate this quality loss, [28] proposed to replace cutoffs with reflection on the constraints boundaries. In addition, [13] extended this reflection to manifolds and introduced log-barrier as a constraint enforcer on manifolds. When saturation is not an issue, projection remains the most widely adopted approach to enforce constraints [10, 22, 34, 35, 48]. While a single projection at the end of inference is sufficient to enforce a constraint [22][10, 48] realized that several projection steps interleaved with the denoising process yield higher quality constraint-satisfying samples. Bringing these advances to diffusion planning, [34] enforces state and action bounds by projecting partially denoised trajectories during inference. Most of these works have also agree on that constraints should mostly be enforced at low noise levels where the sampled predictions are not just random sequences [5, 25, 48]. However, none of these works enforces dynamic feasibility constraints, whose autoregressive nature renders admissible diffusion planning significantly more challenging.
@@ II-B2 Diffusion planning on states
Most closely related to this work is the emerging literature enforcing constraints in diffusion planning with quadratic programming [22], control barrier functions [47], or projections [34]. However, only work [35] considers dynamic constraints, but only test their approach on fully-actuated systems with trivial dynamics.
@@ III Framework
Let us now introduce our framework. We consider a robot of state $s_{t}$ at time $t$ with deterministic discrete-time dynamics of form
@@
where action $a_{t}$ belongs to the admissible action set $\mathcal{A}$ and $\mathcal{S}_{0}$ is the set of initial states in the state space $\mathcal{S}$. We emphasize that dynamics (1) are a black-box, i.e. the equations describing $f$ are not available, but the user can obtain $s_{t+1}$ given $s_{t}$ and $a_{t}$. This setting corresponds to a numerical simulator which is typically available for any robot. We now want to use diffusion policies to generate trajectories of robot (1).
@@ III-B Problem formulation
Equivalently, admissibility requires each state $s_{t+1}$ to belong to the reachable set $\mathcal{R}(s_{t})$ of its predecessor $s_{t}$, which is defined as the set of all feasible next states:
$\mathcal{R}(s_{t}):=\big{\{}f(s_{t},a):\ \text{for all}\ a\in\mathcal{A}\big{\}}.$ (3)
@@ IV-A Reachable sets underapproximation
We assume to know a polytopic underapproximation of the admissible action set $\mathcal{A}$ with vertices $v_{1},...,v_{m}\in\mathcal{A}$ such that $\operatorname{conv}\big{(}v_{1},...,v_{m}\big{)}\subseteq\mathcal{A}$. Then, given a state $s_{t}$, we underapproximate its reachable set with
$\mathcal{C}(s_{t}):=\operatorname{conv}\big{(}f(s_{t},v_{1}),...,f(s_{t},v_{m})\big{)},$ (4)
@@
We illustrate a projection step on Fig. 1 and summarize the whole trajectory projection in Algorithm 1.
@@ IV-C Using action predictions for better state projections
While Algorithms 1 and 2 perform relatively well in low-dimensional state spaces such as the Hopper of Gymnasium [7], these projections are not sufficiently precise to recreate admissible trajectories $\tau$ of the Walker2D and HalfCheetah as $\big{\|}\mathcal{P}(\tau)-\tau\big{\|}>\!\!>1$. This accuracy degradation is most likely caused by the non-convexity of the reachable sets following Remark 1 and the increased action space dimension which enlarges set (4) and keeps the projection far from the reference.
@@ IV-D Using action predictions for admissible state prediction
While projector (11) entirely discards the next state prediction $\tilde{s}_{t+1}$, we can leverage its information to design a smarter projector. We propose to use a feedback correction term $\delta a_{t}$ on the action $\tilde{a}_{t}$ based on the distance between its corresponding next state $\mathcal{P}^{A}$ and the desired next state $\tilde{s}_{t+1}$, as illustrated on Fig. 4. Such a projector $\mathcal{P}^{\text{SA}}$ leverages the combined information of state and action predictions as
@@
To calculate the feedback correction $\delta a_{t}$ we prefer a small neural network $\pi_{\theta}$ to an optimization algorithm because we need $\mathcal{P}^{SA}$ to be differentiable and sufficiently fast to be incorporated into the training of our diffusion models. A model-based approach would not work either due to the black-box nature of $f$. Then, $\delta a_{t}:=\pi_{\theta}\big{(}\tilde{s}_{t+1}-\mathcal{P}^{\text{A}}(s_{t},\tilde{a}_{t})\big{)}$. Given the dataset of trajectories $\mathcal{D}$, we can extract admissible triplets $(s_{t},a_{t},s_{t+1})$, sample a correction term $\delta a_{t}\sim\mathcal{N}(0,\sigma^{2}I)$ corrected next state $s^{\delta}_{t+1}:=f(s_{t},a_{t}+\delta a_{t})$ and train $\pi_{\theta}$ as
@@ V-A Training with projections
1:dataset of trajectories $\mathcal{D}$.
@@
3:     $\sigma\sim\exp\big{(}\mathcal{N}(p_{m},p_{s}^{2})\big{)}$triangleright noise level
@@
To use the reference projector $\mathcal{P}^{\text{ref}}$ of (6) during training, we can obviously use the trajectories from the dataset as references to guide the projection. However, during inference, this dataset might be unavailable, or there might not be a trajectory matching closely the desired one. To solve this issue we leverage the imitation capabilities of diffusion models as illustrated on Fig. 2 where the sampled trajectory matches exactly the desired trajectory. Hence, we use as reference the sampled trajectory before projection, i.e., line 6 of the Inference Algorithm 4 becomes $\tau_{i+1}\leftarrow\mathcal{P}^{\text{ref}}_{\sigma_{i}}\big{(}\tau_{i+1},\tau^{\text{ref}}\leftarrow\tau_{i+1}\big{)}$.
@@ V-B Projection curriculum
where $b$ is a Bernoulli variable taking value $1$ with probability $p(\sigma)$ and 00 otherwise. Thus, projection $\mathcal{P}$ occurs with probability $1-p(\sigma)$. In this transition regime, not projecting all the state transitions of a trajectory breaks the compounding effect and significantly reduces the total projection error as illustrated on Fig. LABEL:fig:_DDAT_scheme.
@@ VI-B Diffusion architecture
The objective of our diffusion models is to predict sequences, thus we chose a transformer backbone instead of the widely used U-Net [20, 32]. We implemented diffusion transformers (DiT) [31] adapted to trajectory predictions by [26] and relying on the notations and best practices of [21]. More details on the diffusion process and our DiT architecture can be found in Appendices C.
To incorporate our trajectory projections into the training of our diffusion policy $D_{\theta}$ as illustrated on Fig. LABEL:fig:_DDAT_scheme, we need our projectors to be differentiable, which is accomplished by using cvxpylayers [1] to solve the convex optimization (10). Indeed, neural network training requires to propagate gradients of the loss $\mathcal{L}=\|\mathcal{P}_{\sigma}(D_{\theta}(\tau+\varepsilon))-\tau\|$ to $D_{\theta}$ and thus have to go through projector $\mathcal{P}_{\sigma}$. Since black-box simulator $f$ is typically not differentiable and is a part of the projections, the gradients are not propagated perfectly through $\mathcal{P}$.
The inference of Algorithm 4 is performed on a batch of trajectories for each initial state $s_{0}$ and is followed by a selection phase to pick the best trajectory according to a desired metric. For most of our robots we select the predicted trajectory satisfying known state constraints for the longest time interval. For instance, on the Hopper, Walker and Unitree this is the time before falling.
@@ VI-C Evaluation metrics
where the minimum is well-defined if $\mathcal{A}$ is compact and $f$ is continuous. When $s_{t+1}$ is reachable from $s_{t}$, the minimum of the norm in (17) is 00 and $f\big{(}s_{t},I\!D(s_{t},s_{t+1})\big{)}=s_{t+1}$. Given a sequence of states $\tilde{\tau}=\big{\{}s_{0},\tilde{s}_{1},...\tilde{s}_{H}\big{\}}\in\mathcal{S}^{H+1}$ we autoregressively apply the inverse dynamics model to find the closest admissible next state to the prediction as detailed in Algorithm 5. Solving optimization problem (17) with sufficient precision to make $I\!D$ a ground-truth model is challenging due to the black-box nature of $f$. We discuss different approaches to solve (17) in Appendix B. Equipped with this inverse dynamics model, we use the following two criteria to quantify the admissibility of trajectories.
@@
Model nomenclature: To distinguish the models evaluated, their name starts with a letter $S$, $A$, or $SA$ depending on whether they predict states, actions or both. The second letter denotes the projector used, i.e., $\mathcal{P}$ (5), $\mathcal{P}^{\text{ref}}$ (6), $\mathcal{P}^{\text{A}}$ (11), or $\mathcal{P}^{\text{SA}}$ (12). The projector can take two subscripts: $\sigma$ if a projection curriculum is used ($\sigma_{min}=0.0021$, $\sigma_{max}=0.2$ unless otherwise specified) and $I$ if the projector is only used at inference, it is otherwise assumed to be also used during training. Finally, conditioning of models on a variable $v$ is denoted by $(\cdot|v)$.
@@ VII-B Projection curriculum during inference
Table II and Figures 7(a) and 7(d) show no significant difference in terms of admissibility based on the projection curriculum. However, the quality of the trajectories is significantly reduced by projections at high noise level as shown on Table II and Fig. 5(a) and 6. This answers Q2.
Figure 6: Quadcopter trajectories generated by the $SA\mathcal{P}_{\sigma}^{\text{SA}}$ model with projections starting either pre-inference, or mid-inference or post-inference. The desired behavior exhibited on Fig. 8(c) slaloms between the obstacles. None of the trajectories generated by the pre-inference model pass the obstacles, whereas $81\%$ and $79\%$ of the trajectories generated by the other two projection curricula pass the slalom. Hence, projections at the beginning of inference strongly reduce the quality of the samples.

### v2

# {textblock}

@@ Abstract
Diffusion models excel at creating images and videos thanks to their multimodal generative capabilities. These same capabilities have made diffusion models increasingly popular in robotics research, where they are extensively used for generating robot motion. However, the stochastic nature of diffusion models is fundamentally at odds with the precise dynamical equations describing the feasible motion of robots. Hence, generating dynamically admissible robot trajectories is a challenge for diffusion models. To alleviate this issue, we introduce DDAT: Diffusion policies for Dynamically Admissible Trajectories to generate admissible trajectories of black-box robotic systems using diffusion models. To generate such trajectories, our diffusion policies project their predictions onto a dynamically admissible manifold during both training and inference to align the objective of the denoiser neural network with the dynamical admissibility constraint. Due to the auto-regressive nature of such projections as well as the black-box nature of robot dynamics, trajectory projections are challenging. We thus enforce admissibility by iteratively sampling a polytopic under-approximation of the reachable set of a state onto which we project its predicted successor, before iterating this process with the projected successor. By producing accurate trajectories, this projection eliminates the need for diffusion models to continually replan, enabling one-shot long-horizon trajectory planning. We demonstrate that our proposed framework generates higher quality dynamically admissible robot trajectories through extensive simulations on a quadcopter and various MuJoCo environments, along with real-world experiments on a Unitree GO1 and GO2. Code available on our website: https://iconlab.negarmehr.com/DDAT/.
@@ I Introduction
To address the dynamic feasibility issue in diffusion planning from offline data, we introduce Diffusion policies for Dynamically Admissible Trajectories (DDAT). These models use a chronological trajectory projector capable of transforming a generated sequence of states into an admissible trajectory. On the contrary to work [22], and inspired by [10, 16, 50], the projections in DDAT occurs throughout and after inference instead of only after inference. Differing from works [10, 16, 50], these projections are also implemented during training of the diffusion models to align their training objectives with their inference phase, as suggested in work [5]. Furthermore, works [5, 10, 16, 50] only concern themselves with a single constraint, which is significantly less challenging to enforce than dynamical admissibility due to its autoregressive nature. Another advantage of generating feasible trajectories is that we do not need to sample and denoise a large batch of trajectories hoping to find an admissible one, which provides significant computational gain compared to methods like [8]. Only works [31, 36] studied dynamical feasibility in diffusion planning, but both works focus solely on linear systems.
Our main contributions in this work are as follows.
@@
    We generate dynamically admissible trajectories with diffusion models for black-box systems.
@@
The remainder of this work is organized as follows. In Section II, we provide a survey of related works. In Section III, we introduce our problem formulation along with our framework. In Section IV, we define several projectors to enforce the dynamical admissibility of trajectories. In Section V, we incorporate these projectors into diffusion models. In Section VI, we implement our proposed Diffusion policies for Dynamically Admissible Trajectories. We present our simulation results on a quadcopter, the Gym Hopper, Walker and HalfCheetah, and the Unitree GO1 and GO2 quadrupeds in Section VII, while Section VIII demonstrates the hardware implementation of DDAT on a Unitree GO1 and a GO2.
@@ II-A2 Hard constraints
Contrasting the works presented above, we are interested in hard constraints to guarantee their satisfaction. The first hard constraints imposed on diffusion models were $[0,255]$ pixel cutoffs, which led to saturated images [28]. To mitigate this quality loss, [28] proposed to replace cutoffs with reflection on the constraints boundaries. In addition, [13] extended this reflection to manifolds and introduced log-barrier as a constraint enforcer on manifolds. When saturation is not an issue, projection remains the most widely adopted approach to enforce constraints [10, 22, 35, 36, 50]. While a single projection at the end of inference is sufficient to enforce a constraint [22], works [10, 50] realized that several projection steps interleaved with the denoising process yield higher quality constraint-satisfying samples. Bringing these advances to diffusion planning, [35] enforces state and action bounds by projecting partially denoised trajectories during inference. Most of these works also agree that constraints should mostly be enforced at low noise levels where the sampled predictions are not just random sequences [5, 25, 50]. However, none of these works enforces dynamic feasibility constraints, whose autoregressive nature renders admissible diffusion planning significantly more challenging.
@@ II-B2 Diffusion planning on states
Most closely related to this work is the emerging literature enforcing constraints in diffusion planning with quadratic programming [22], control barrier functions [49], or projections [35]. Only works [31, 36] consider dynamic constraints, but only test their approaches on linear fully-actuated systems.
@@ III Framework
Let us now introduce our framework. We consider a robot of state $s_{t}$ at time $t$ with black-box deterministic discrete-time dynamics of form
@@
where action $a_{t}$ belongs to the admissible action set $\mathcal{A}\subset\mathbb{R}^{m}$ and $\mathcal{S}_{0}$ is the set of initial states in the state space $\mathcal{S}\subseteq\mathbb{R}^{n}$. We emphasize that dynamics (1) are a black-box, i.e. the equations describing $f$ are not available, but the user can obtain $s_{t+1}$ given $s_{t}$ and $a_{t}$. This setting corresponds to a numerical simulator which is typically available for any robot. We now want to use diffusion policies to generate trajectories of robot (1).
@@ III-B Problem formulation
Equivalently, admissibility requires each state $s_{t+1}$ to belong to the reachable set of its predecessor $s_{t}$.
Definition 2.
The reachable set $\mathcal{R}(s_{t})$ of a state $s_{t}$ is the set of all feasible next states:
$\mathcal{R}(s_{t}):=\big{\{}f(s_{t},a):\ \emph{\text{for all}}\ a\in\mathcal{A}\big{\}}.$ (3)
@@ IV-A Reachable sets underapproximation
We assume to know a polytopic underapproximation of the admissible action set $\mathcal{A}$ with vertices $v_{1},...,v_{q}\in\mathcal{A}$ such that $\operatorname{conv}\big{(}v_{1},...,v_{q}\big{)}\subseteq\mathcal{A}$. This assumption is typically verified as $\mathcal{A}$ is often a known hyperrectangle of dimension $m$ delimited by $q=2^{m}$ vertices [7]. Then, given a state $s_{t}$, we underapproximate its reachable set with
$\mathcal{C}(s_{t}):=\operatorname{conv}\big{(}f(s_{t},v_{1}),...,f(s_{t},v_{q})\big{)},$ (4)
@@
We illustrate a projection step in Fig. 1 and summarize the whole trajectory projection in Algorithm 1.
Remark 1.
@@ IV-C Using action predictions for better state projections
While Algorithms 1 and 2 perform relatively well in low-dimensional state spaces such as the Hopper of Gymnasium [7], these projections are not sufficiently precise to recreate admissible trajectories $\tau$ of the Walker2D and HalfCheetah as $\big{\|}\mathcal{P}(\tau)-\tau\big{\|}>\!\!>0$. This accuracy degradation is most likely caused by the non-convexity of the reachable sets following Remark 1 and the increased action space dimension which enlarges set (4) and keeps the projection far from the reference.
@@ IV-D Using action predictions for admissible state prediction
While projector (11) entirely discards the next state prediction $\tilde{s}_{t+1}$, we can leverage its information to design a smarter projector. We propose to use a feedback correction term $\delta a_{t}$ on the action $\tilde{a}_{t}$ based on the distance between its corresponding next state $\mathcal{P}^{A}$ and the desired next state $\tilde{s}_{t+1}$, as illustrated in Fig. 4. Such a projector $\mathcal{P}^{\text{SA}}$ leverages the combined information of state and action predictions as
@@
To calculate the feedback correction $\delta a_{t}$ we prefer a small neural network $\pi_{\theta}$ to an optimization algorithm because we need $\mathcal{P}^{SA}$ to be differentiable and sufficiently fast to be incorporated into the training of our diffusion models. A model-based approach would not work either due to the black-box nature of $f$. Then, $\delta a_{t}:=\pi_{\theta}\big{(}\tilde{s}_{t+1}-\mathcal{P}^{\text{A}}(s_{t},\tilde{a}_{t})\big{)}$. Given the dataset of trajectories $\mathcal{D}$, we can extract admissible triplets $(s_{t},a_{t},s_{t+1})$, sample a correction term $\delta a_{t}\sim\mathcal{N}(0,\sigma^{2}I)$, correct next state $s^{\delta}_{t+1}:=f(s_{t},a_{t}+\delta a_{t})$, and train $\pi_{\theta}$ as
@@ V-A Training with projections
1:dataset of trajectories $\mathcal{D}$, projector $\mathcal{P}_{\sigma}$ of (5), (6), (11), or (12).
@@
3:     $\sigma\sim\exp\big{(}\mathcal{N}(p_{m},p_{s}^{2})\big{)}$triangleright noise level, Appendix C.
@@
To use the reference projector $\mathcal{P}^{\text{ref}}$ of (6) during training, we can obviously use the trajectories from the dataset as references to guide the projection. However, during inference, this dataset might be unavailable, or there might not be a trajectory matching closely the desired one. To solve this issue we leverage the imitation capabilities of diffusion models as illustrated in Fig. 2 where the sampled trajectory matches exactly the desired trajectory. Hence, we use as reference the sampled trajectory before projection, i.e., line 6 of the Inference Algorithm 4 becomes $\tau_{i+1}\leftarrow\mathcal{P}^{\text{ref}}_{\sigma_{i}}\big{(}\tau_{i+1},\tau^{\text{ref}}\leftarrow\tau_{i+1}\big{)}$.
@@ V-B Projection curriculum
where $b$ is a Bernoulli variable taking value $1$ with probability $p(\sigma)$ and 00 otherwise. Thus, projection $\mathcal{P}$ occurs with probability $1-p(\sigma)$. In this transition regime, not projecting all the state transitions of a trajectory breaks the compounding effect and significantly reduces the total projection error as illustrated in Fig. LABEL:fig:_DDAT_scheme.
@@ VI-B Diffusion architecture
The objective of our diffusion models is to predict sequences, thus we chose a transformer backbone instead of the widely used U-Net [20, 33]. We implemented diffusion transformers (DiT) [32] adapted to trajectory predictions by [26] and relying on the notations and best practices of [21]. More details on the diffusion process and our DiT architecture can be found in Appendix C.
To incorporate our trajectory projections into the training of our diffusion policy $D_{\theta}$ as illustrated in Fig. LABEL:fig:_DDAT_scheme, we need our projectors to be differentiable, which is accomplished by using cvxpylayers [1] to solve the convex optimization (10). Indeed, neural network training requires to propagate gradients of the loss $\mathcal{L}=\|\mathcal{P}_{\sigma}(D_{\theta}(\tau+\varepsilon))-\tau\|$ to $D_{\theta}$ and thus have to go through projector $\mathcal{P}_{\sigma}$. Since the black-box simulator $f$ is typically not differentiable and is a part of the projections, the gradients are not propagated perfectly through $\mathcal{P}$.
The inference of Algorithm 4 is performed on a batch of trajectories for each initial state $s_{0}$ and is followed by a selection phase to pick the best trajectory according to a desired metric. For most of our robots we select the predicted trajectory satisfying known state constraints for the longest time interval. For instance, on the Hopper, Walker, and Unitree this is the time before falling.
@@ VI-C Evaluation metrics
where the minimum is well-defined if $\mathcal{A}$ is compact and $f$ is continuous. When $s_{t+1}$ is reachable from $s_{t}$, the minimum of the norm in (17) is 00 and $f\big{(}s_{t},I\!D(s_{t},s_{t+1})\big{)}=s_{t+1}$. Given a sequence of states $\tilde{\tau}=\big{\{}s_{0},\tilde{s}_{1},...\tilde{s}_{H}\big{\}}\in\mathcal{S}^{H+1}$ we autoregressively apply the inverse dynamics model to find the closest admissible next state to the prediction as detailed in Algorithm 5 located in Appendix B. Solving optimization problem (17) with sufficient precision to make $I\!D$ a ground-truth model is challenging due to the black-box nature of $f$. We discuss different approaches to solve (17) in Appendix B. Equipped with this inverse dynamics model, we use the following two criteria to quantify the admissibility of trajectories.
@@
Model nomenclature: To distinguish the models evaluated, their name starts with a letter $S$, $A$, or $SA$ depending on whether they predict states, actions or both. The second part denotes the projector used, i.e., $\mathcal{P}$ (5), $\mathcal{P}^{\text{ref}}$ (6), $\mathcal{P}^{\text{A}}$ (11), or $\mathcal{P}^{\text{SA}}$ (12). The projector can take two subscripts: $\sigma$ if a projection curriculum is used and $I$ if the projector is only used at inference, it is otherwise assumed to be also used during training. Finally, conditioning of models on a variable $v$ is denoted by $(\cdot|v)$.
@@ VII-B Projection curriculum during inference
Table II and Figures 7(a) and 7(d) show no significant difference in terms of admissibility based on the projection curriculum. However, the quality of the trajectories is significantly reduced by projections at high noise level as shown in Table II and Fig. 5(a) and 6. This answers Q2.
Figure 6: Quadcopter trajectories generated by the $SA\mathcal{P}_{\sigma}^{\text{SA}}$ model with projections starting either pre-inference, or mid-inference or post-inference. The desired behavior exhibited in Fig. 8(c) slaloms between the obstacles. None of the trajectories generated by the pre-inference model pass the obstacles, whereas $81\%$ and $79\%$ of the trajectories generated by the other two projection curricula pass the slalom. Hence, projections at the beginning of inference strongly reduce the quality of the samples.